OpenStudy (milo123):

MEDAL!!! Matt looks at the architectural plan of a four-walled room in which the walls meet each other at right angles. The length of one wall in the plan is 19 inches. The length of the diagonal of the floor of the room in the plan is approximately 20.62 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work.

OpenStudy (milo123):

1. the walls meet each other at right angles. (square all angles are 90 deg) 2. a2+c2=b2 right angle triangle hypotenuse theorem with a = 19 and b = 20.62, find c if c = a then sides are equal so it is a square

OpenStudy (milo123):

how can i make this sound better/make more sense....if this is correct? @jigglypuff314 @Jaynator495

OpenStudy (greison14):

hey do u still need help or no

OpenStudy (halloweeninheaven):

1. the walls meet each other at right angles. (square all angles are 90 deg) 2. a2+c2=b2 right angle triangle hypotenuse theorem with a = 19 and b = 20.62, find c if c = a then sides are equal

OpenStudy (halloweeninheaven):

Pythagoras Theorem*

OpenStudy (milo123):

yes i do need help:) but only if you really know what to do and can help lol..^he just copied that from the internet

OpenStudy (radar):

The diagonal is the hypotenuse, so if:\[19^{2}+ 19^{2}= 20.62^{2}\] If it is an equality, then the room is square.

OpenStudy (radar):

I hope that made sense, and good luck with your studies.

OpenStudy (milo123):

wait...

OpenStudy (radar):

Yessssss.

OpenStudy (milo123):

It is about 19 inches by 8 inches. so its not a square?

OpenStudy (milo123):

or i'm wrong lol

OpenStudy (milo123):

19^2+19^2=722 square root of 722 is 28.87

OpenStudy (radar):

Well lets us verify. 19 squared is 361 361+361= 722 now what is 20.62 squared??? 425 not even close. The room is not square.

OpenStudy (milo123):

425.18 Square root of 425.18 = 20.61 .....

OpenStudy (radar):